# From Models to Network Topologies: A Topology Inference Attack in Decentralized Federated Learning

### v2

@@ 6.3 Attack Performance
After establishing the attack metrics, the proposed topology inference attacks were evaluated on five datasets, encompassing four attack scenarios and different topologies. The results in Table 1 demonstrate that the proposed methods, EDGEPRE and INFERGAT, are highly effective, achieving the best performance across all four attack scenarios and five datasets.
Table 1: Averaged Topology inference performance (F1-Score) under four attack scenarios across five datasets.
Scenario Algorithm CIFAR10 FMNIST ImageNet10 MNIST SVHN
SC1 Logistic 0.704 0.773 0.752 0.795 0.788
RF 0.844 0.845 0.879 0.844 0.851
SVM 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.851 0.875 0.840 0.876 0.874
SC2 Logistic 0.704 0.771 0.752 0.794 0.787
RF 0.844 0.845 0.879 0.844 0.851
SVC 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.850 0.875 0.843 0.876 0.872
SC3 GMM 0.723 0.443 0.482 0.669 0.679
Kmeans 0.727 0.715 0.618 0.724 0.734
Spectral 0.597 0.638 0.591 0.607 0.648
INFERGAT 0.800 0.788 0.733 0.799 0.802
SC4 GMM 0.576 0.662 0.685 0.670 0.688
Kmeans 0.730 0.776 0.741 0.785 0.795
Spectral 0.683 0.645 0.636 0.637 0.627
INFERGAT 0.797 0.810 0.743 0.807 0.812
It can also be observed that from SC1 to SC4, the attacker’s accessible knowledge gradually decreases, leading to a corresponding decline in attack performance. In SC1, the inferred topologies achieve an F-score exceeding 85% on all five datasets, whereas in SC4, the score drops to approximately 80%. This indicates that prior knowledge has a significant impact on the success rate of topology inference. Comparing SC1 and SC2, the difference in performance is relatively small, suggesting that combining data and model knowledge offers limited additional benefit over model-only knowledge. This implies that the model itself already encodes sufficient information to support effective topology inference.
@@
A. Effect of Topology Density. To investigate the influence of topological characteristics on topology inference attacks, Figure 2 presents the F1-Scores achieved by each attack scenario under five in the CIFAR10 dataset. Topology density represents the ratio of the actual number of edges to the maximum possible number of edges, calculated as $\frac{2E}{N(N-1)}$. This metric reflects the overall connectivity or sparsity of a graph.
As shown in Figure 2, across four different attack scenarios, a consistent trend was observed: as the graph’s density increases, the attack success rate of the inference models (EDGEPRE and INFERGAT) tends to decrease. This is primarily because denser topologies lead to increased information fusion, resulting in higher similarity among local models. In highly dense networks, the models become increasingly homogeneous, making it more difficult for the inference algorithms to distinguish whether a connection exists between nodes. In contrast, sparse topologies exhibit greater heterogeneity among nodes despite being limited by data sparsity and class imbalance. This heterogeneity simplifies the inference task, allowing the attack models to make more accurate predictions.
@@
Figure 2: Impact of topology density on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
B. Effect of Network Size. To evaluate the scalability of the proposed inference algorithms under different attack scenarios, Figure 3 illustrates the performance variation across four attack scenarios as the network size increases. The results show a consistent upward trend in attack effectiveness with increasing numbers of nodes, demonstrating the strong scalability of the proposed methods. On one hand, the algorithms accurately infer the underlying topology of DFL networks across varying levels of density, indicating their robustness in handling data imbalance challenges. On the other hand, even in large-scale networks with up to 161 nodes, the inference models maintain strong and stable performance, further validating the stability and scalability of the proposed topology inference approaches.
@@
Figure 3: Impact of network size on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
@@ 7 Mitigation
Figure 4: Impact of local training epochs, data augmentation, and data heterogeneity on the F1-Score of topology inference attacks in the CIFAR10 dataset.
@@ 7.1 Reduce Overfitting
As illustrated in Figure 4, each bar represents a specific training configuration: gray bars denote IID data distributions without augmentation, blue bars correspond to non-IID data distributions (Dirichlet $\alpha=0.1$), and orange bars indicate the use of data augmentation under IID settings.
Consistent with findings in prior inference attack literature, the success of topology inference attacks is closely tied to the degree of overfitting in the target model. To validate this, experiments were conducted on the CIFAR-10 dataset under four attack scenarios. By increasing the number of local training epochs from 3 to 10, the level of overfitting was elevated. Results show that, across all examined topologies, this change consistently improves the F1-Scores of attacks. This observation confirms that overfitting exacerbates the risk of topological information leakage in DFL systems.
To counteract overfitting, data augmentation techniques were employed, specifically random cropping with 4-pixel padding. The impact is evident in the substantial performance decline observed for the orange bars across all scenarios and topologies. This demonstrates that data augmentation not only enhances the generalization ability of local models but also effectively mitigates the success of topology inference attacks.
@@ 7.2 Data Heterogeneity
Overall, these findings imply that reducing overfitting through fewer local training epochs or data augmentation and increasing data heterogeneity are promising directions for mitigating topology inference risks in decentralized federated learning.
@@ 8 Conclusion and Future Work
This work introduces a novel topology inference attack against DFL, exposing critical vulnerabilities related to privacy and information leakage. By analyzing local models, attackers can accurately infer the overlay topology, one of DFL’s most sensitive assets, highlighting the system’s susceptibility to privacy breaches. The study explores various attack scenarios and develops tailored metrics, models, and algorithms, with experiments confirming the feasibility and effectiveness of these attacks. Furthermore, network size, density, model overfitting, and data heterogeneity significantly influence attack success. Mitigation strategies, such as data augmentation and increasing data heterogeneity, can enhance model generalization and reduce the risk of data leakage.

### v3

@@ 6.3 Attack Performance
After establishing the attack metrics, the proposed topology inference attacks are evaluated on six datasets under four scenarios. As shown in Table 1, EDGEPRE dominates in SC1–SC2, while INFERGAT ranks best in SC3–SC4 across all datasets.
Performance decreases as the attacker’s prior knowledge is reduced from SC1 to SC4. In SC1, EDGEPRE attains F1 of at least $\approx$0.85 on every dataset (with high AUC as well). Moving to SC2 yields only a modest drop (typically a few points), indicating that model parameters alone already encode most of the topology signal. The larger degradation appears when supervision is removed: in SC3, INFERGAT remains strongest but F1 becomes dataset-dependent (e.g., $\approx$0.89 on CIFAR10 vs. $\approx$0.66–0.77 on MNIST and Malware). Under the weakest prior (SC4), INFERGAT still outperforms other unsupervised baselines, achieving F1 around 0.60–0.67 with AUC $\approx$0.76–0.80. Overall, these trends confirm that stronger prior/model knowledge markedly improves topology inference, whereas limited priors and no labels (SC3 and SC4) lead to the main loss in accuracy.
@@
A. Effect of Topology Density. To investigate the influence of topological characteristics on topology inference attacks, Figure 2 reports the F1-Scores for four attack scenarios on the CIFAR10 dataset. Topology density, defined as $\frac{2E}{N(N-1)}$, reflects overall connectivity.
Overall, the four scenarios exhibit relatively stable performance across the full range of densities. For SC3 and SC4, which represent unsupervised attacks, denser topologies provide richer edge information that supports the recovery of the topology, leading to gradually improved performance as connectivity increases. In contrast, SC1 and SC2 achieve comparatively high scores across all densities, reflecting their ability to exploit both sparse and dense settings. The shaded areas further indicate that variability is higher at very low densities, where limited connections introduce uncertainty, while the results become stable once sufficient connectivity is reached.
@@
Figure 2: Impact of topology density on the F1-scores achieved by each attack scenario on the CIFAR10 dataset.
B. Effect of Network Size. To evaluate the scalability of the inference algorithms under different attack scenarios, Figure 3 illustrates the variation in F1-Scores as the network size increases. SC1 and SC2 remain relatively stable across different numbers of nodes, indicating robustness to changes in network size. In contrast, SC3 and SC4 exhibit a clear decline as the number of nodes grows. This behavior is explained by the fact that larger networks typically correspond to lower densities, which reduce the amount of effective edge information available for unsupervised inference. Consequently, the performance of SC3 and SC4 degrades with increasing network size, whereas the supervised scenarios (SC1 and SC2) maintain more stable effectiveness.
@@
Figure 3: Impact of network size on the F1-scores achieved by each attack scenario on the CIFAR10 dataset.
@@ 7 Mitigation
Figure 4: Impact of local training epochs and data heterogeneity on the F1-Score of topology inference attacks in the CIFAR10 dataset.
@@ 7.1 Reduce Overfitting
As illustrated in Figure 4, each bar represents a specific training configuration: gray bars denote IID data distributions without augmentation and blue bars correspond to non-IID data distributions (Dirichlet $\alpha=0.1$).
Consistent with findings in prior inference attack literature, the success of topology inference attacks is closely tied to the degree of overfitting in the target model. To validate this, experiments were conducted on the CIFAR10 dataset under four attack scenarios. By increasing the number of local training epochs from 3 to 10, the level of overfitting was elevated. Results show that, across all examined topologies, this change consistently improves the F1-Scores of attacks. This observation confirms that overfitting exacerbates the risk of topological information leakage in DFL systems.
@@ 7.2 Data Heterogeneity
7.3 Differential Privacy
Differential Privacy (DP) is a common mitigation for privacy in FL [2]. DP is integrated into DFL parameter sharing and its effect on topology inference is assessed. On CIFAR10 (ER_0.5) under SC1, attack F1 drops from $0.889\rightarrow 0.687$ (10 nodes) and $0.824\rightarrow 0.637$ (20 nodes), approximately $23\%$ reductions, indicating effective mitigation.
Overall, these findings imply that reducing overfitting through fewer local training epochs and increasing data heterogeneity are promising directions for mitigating topology inference risks in DFL.
@@ 8 Conclusion and Future Work
This work introduces a novel topology inference attack against DFL, exposing critical vulnerabilities related to privacy and information leakage. By analyzing local models, attackers can accurately infer the overlay topology, one of DFL’s most sensitive assets, highlighting the system’s susceptibility to privacy breaches. The study explores various attack scenarios and develops tailored metrics, models, and algorithms, with experiments confirming the feasibility and effectiveness of these attacks. Furthermore, network size, density, model overfitting, and data heterogeneity significantly influence attack success. Mitigation strategies, such as increasing data heterogeneity, can enhance model generalization and reduce the risk of data leakage.
@@
Acknowledgment
This work has been partially supported by (a) the Swiss Federal Office for Defense Procurement (armasuisse) with the CyberDFL project (CYD-C-2020003) and (b) the University of Zürich UZH.
Appendix A appendix
A.1 Topology Information
This work utilizes two categories of topologies: synthetic and real-world. The synthetic topologies include ring, star, and three Erdős-Rényi (ER) graphs with connection probabilities p=p=0.3, 0.5, and 0.7. The real-world topologies consist of 27 network topologies collected from the two public network repositories: SNDlib [^2] Available in: https://sndlib.put.poznan.pl/home.action and DEFO [^3] Available in: https://sites.uclouvain.be/defo/, which reflect realistic communication and infrastructure networks.
To characterize the structure of different graphs, this work reports the following topological statistics:
  • Number of Nodes: The total number of distinct vertices in the graph, denoted as $N$.
  • Number of Edges: The total number of edges connecting node pairs, denoted as $E$.
  • Average Degree: The mean number of edges connected to each node, calculated as $\frac{2E}{N}$.
  • Density: The ratio of the number of actual edges to the maximum possible number of edges, given by $\frac{2E}{N(N-1)}$. This metric reflects the graph’s overall connectivity or sparsity.
A.1.1 Synthetic Topology Characterize Statistics
To evaluate the robustness and generalizability of the proposed method under various topology conditions, this work considers a set of synthetic topologies ranging from sparse to dense:
  • STAR graphs represent extremely sparse topologies, where most nodes are only connected to a single central hub. These graphs have low average degree and density, reflecting minimal connectivity.
  • RING graphs are sparse structures with uniform and low node degrees. While slightly denser than STAR graphs, they remain sparse and exhibit limited path diversity.
  • ER graphs with varying edge probabilities ($p=0.3$, $0.5$, and $0.7$) capture a continuum of connectivity levels. As $p$ increases, the average degree and edge density grows, gradually transitioning the graph from a sparse to a dense structure.
This selection enables controlled evaluation across different topological regimes, ensuring that performance is not biased toward a specific graph sparsity or structure. The characterize statistics of synthetic topologies are shown in Table 2
Table 2: Topological Statistics of Different Synthetic Topology
Topology Number of Nodes Number of Edges Average Degree Density
ER($p=0.3$) 10 16 3.20 0.36
ER($p=0.3$) 30 115 7.67 0.26
ER($p=0.3$) 20 50 5.00 0.26
ER($p=0.5$) 10 23 4.60 0.51
ER($p=0.5$) 30 202 13.47 0.46
ER($p=0.5$) 20 87 8.70 0.46
ER($p=0.7$) 20 132 13.20 0.69
ER($p=0.7$) 30 298 19.87 0.69
ER($p=0.7$) 10 28 5.60 0.62
RING 10 10 2.00 0.22
RING 20 20 2.00 0.11
RING 30 30 2.00 0.07
STAR 10 9 1.80 0.20
STAR 20 19 1.90 0.10
STAR 30 29 1.93 0.07
A.1.2 Real-world Topology Characterize Statistics
To evaluate the generalizability and robustness of the proposed approach across diverse topological structures, experiments were conducted on a wide range of real-world network topologies. These networks vary significantly in size, ranging from small-scale graphs with only 10 nodes to large-scale structures with up to 87 nodes. The topologies are sourced from realistic communication infrastructures, including national research and education networks (NRENs) and Internet service provider (ISP) backbone networks, which reflect actual deployment scenarios.
A statistical overview of these networks is provided in Table 3. Compared to synthetic graphs, real-world networks exhibit higher structural variability in terms of node degree, connectivity, and scale. Despite this diversity, they are generally characterized by sparse connectivity, as indicated by low edge densities and moderate average degrees. This sparsity is a common property of real communication networks and introduces specific challenges to graph data.
Table 3: Topological statistics of real-world and synthetic networks
Topology Number of Nodes Number of Edges Average Degree Density
ABILENE 12 15 2.50 0.11
ATLANTA 15 22 2.93 0.10
COST266 37 57 3.08 0.04
DFN-GWIN 11 47 8.55 0.43
DI-YUAN 11 42 7.64 0.38
FRANCE 25 45 3.60 0.08
GEANT 22 72 6.55 0.16
GERMANY50 50 88 3.52 0.04
GIUL39 39 172 8.82 0.12
INDIA35 35 80 4.57 0.07
JANOS-US 26 84 6.46 0.13
JANOS-US-CA 39 122 6.26 0.08
NEWYORK 16 49 6.13 0.20
NOBEL-EU 28 41 2.93 0.05
NOBEL-GERMANY 17 26 3.06 0.10
NOBEL-US 14 21 3.00 0.12
NORWAY 27 51 3.78 0.07
PDH 11 34 6.18 0.31
PIORO40 40 89 4.45 0.06
POLSKA 12 18 3.00 0.14
RF1755 87 322 7.40 0.04
RF3967 79 294 7.44 0.05
SUN 27 102 7.56 0.15
SYNTH50 50 276 11.04 0.11
TA1 24 51 4.25 0.09
TA2 65 108 3.32 0.03
ZIB54 54 80 2.96 0.03
A.2 Attack Metrics Distribution
This experiment comprehensively demonstrates the distribution of attack metrics in the edge and non-edge groups across CIFAR10 (Figure 5), MNIST(Figure 6), FMNIST(Figure 7), SVHN(Figure 8), ImageNet10 (Figure 9), and Malware (Figure 10) datasets in various topologies. It can be observed that, in most cases, the distribution of relative loss is more concentrated than that of relative entropy and sensitivity. Additionally, relative loss exhibits less overlap between the edge group and the non-edge group. Therefore, this work selects relative loss as the attack metric when the attacker has access to both local data and the local model, i.e., attack scenarios 1 and 3.
Figure 5: Normalized distributions of attack metrics in various topology DFL systems trained on the CIFAR10 dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Figure 6: Normalized distributions of attack metrics in various topology DFL systems trained on the MNIST dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Figure 7: Normalized distributions of attack metrics in various topology DFL systems trained on the FMNIST dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Figure 8: Normalized distributions of attack metrics in various topology DFL systems trained on the SVHN dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Figure 9: Normalized distributions of attack metrics in various topology DFL systems trained on the ImageNet10 dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Figure 10: Normalized distributions of attack metrics in various topology DFL systems trained on the Malware dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection. In each subfigure, the attack metrics are respectively: Relative Loss, Relative Entropy, Relative Sensitivity, Cosine Similarity, Euclidean Similarity, and Curvature Divergence.
Regarding similarity-based metrics, the cosine similarity between connected models is distinctly higher than that between non-connected models, displaying no overlap between the distributions. On the other hand, the curvature divergence and Euclidean similarity show a considerably more significant degree of overlap. Thus, the cosine similarity has been adopted as the attack metric in scenarios where only local models are available (attack scenarios 2 and 4).